In the right triangle find the length of the side not given. Give an exact answer and an approximation to three decimal places. A=2 and B=4 thanks.
Answer by Jk22(389)

2 answer :
1) 4^2+2^2=20=C1^2
2) 4^2=2^2+10=2^2+C2^2
C1=Sqrt(20) : 4.472
20....|4,472
_______________
16....|
_______________
40.0..|84x4=336
336...|
________________
640.0.|887x7=6209
6209..|
_________________
1910.0|8942x2
C2=Sqrt(10) : 3.163
10.....|3,163
9......|
-----------
10.0...|61x1=61
61.....|
-----------
390.0..|626x6=3756
3756...|
-----------
2440.0 |6322x2=..